The IFO alignment is somewhat OK, in that the green and IR beams are flashing in the arms, and the return beams are overlapping at the BS. However the beams appear to be not centered on any of the optics at the moment. They are all displaced in yaw by ~0.5 to 1 cm or so in various directions.
From this we have decided that we need to step back and reattack the IFO alignment from square one. Here is our current suggested procedure:
Once this is done, we will need to check the following:
We've decided to stop for the night, get a good nights rest, and attack all of this tomorrow morning.